2 inches of heavy, wet snow are equivalent to an inch of rain. Estimate the water content of 5 inches of heavy, wet snow. The water content of 5 inches of snow is _____ inches
Answer by jim_thompson5910(35256)

2/1 = 5/x

2*x =5*1

2x = 5

x = 5/2

x = 2.5

The water content of 5 inches of snow is __2.5___ inches